I just need part B answered, thanks!

In a model AC generator, a 504 turn rectangular co
Answer #1

Ans

Emf= -N\frac{\partial \phi }{\partial t}

Where N is no of turn.

\phi = B.A =BA cos\theta

( \theta is the angle between magnetic field B and area vector A )

is the magnetic flux.

t is time.

Also using equation of circular motion

\theta = \theta ^{_{_{_{0}}}} + \omega t

therefore

Emf =-N\frac{\partial BAcos(\omega t +\theta _{0})}{\partial t}

            = Emf =NBA\omega sin(\omega t+\theta _{0})

At t=0, Emf=0;

i.e    sin(\theta _{0}) = 0

or   \theta _{0} = 0

therefore Emf =NBA\omega sin(\omega t)

                            ( \omega = 2\pi *120/60 =4\pi in rad per sec)

                            = 504*0.61*8*26*10-4*4*3.14*sin(4*\pi2/32)

                              =75.78 volt